OpenStudy (raffle_snaffle):

Find the following using the given vectors: a=<1, 1, -2>, b=<3, -2 , 1> a) a*b b)Find the angle between vectors a and b. c) Find the unit vector parallel to b d)axb I found a, b, and d. I am not sure how to solve part (c)

OpenStudy (psymon):

I THINK we need to find the projection of u onto v here. So we're looking for vector components. The projection of u onto v is \[(\frac{ u\dot~v }{ ||v||^{2} })*v\]

OpenStudy (psymon):

Because when we find vector components, one component is parallel and the otheris orthogonal.

OpenStudy (raffle_snaffle):

we have not learned about projections yet...

OpenStudy (psymon):

Oh O.o I see, lol. Whoops!

OpenStudy (raffle_snaffle):

well, we won't be going over projections since we ran out of time. Lol

OpenStudy (raffle_snaffle):

Is that the formula?

OpenStudy (raffle_snaffle):

let me read up on this and i will get back to you guys.

OpenStudy (psymon):

Yeah, to find \[proj _{v}u\] projection of u onto v, which can bewritten as just \[w _{1}\]The second component, the one orthogonal to the vector is: \[u-w _{1}\] And alright, go for it, lol.

OpenStudy (raffle_snaffle):

so what is the difference between the scalar projection and vector project? I see the formulas are quite a bit different

OpenStudy (raffle_snaffle):

Oh nvm I just read it. Lol

OpenStudy (raffle_snaffle):

Okay I got this

OpenStudy (psymon):

Lol, awesome. Yeah, otherwise you'd justinvent yourown vector, but the components let you find a parallel and an orthogonal component. Oh, hey, you're at the beginning of calc 3, right?

OpenStudy (raffle_snaffle):

do you know how to invent your own vector?

OpenStudy (raffle_snaffle):

I am at the end of calc 3

OpenStudy (psymon):

Not really, lol. But I wouldthink that all you'd need to do is shift each coordinate up 1, haha. But that's cheating.

OpenStudy (raffle_snaffle):

we were introduced diff eqs and vectors

OpenStudy (psymon):

Ah. Yeah, I noticed you're doing stuff I didin calc 2. But I was curious because I have the textbook for calc 3 and am just starting it. I wanna keep up with ya, it'd be awesome to have someone studying the same stuff xD

OpenStudy (raffle_snaffle):

yeah for sure. i am starting calc 4 fall term

OpenStudy (psymon):

Ah. I don't think we even have a calc 4 here. Calc 3 is highest. But nah, if we have the same material I'd wanna keep up and learn it as you do.

OpenStudy (raffle_snaffle):

calc 2 was by far the most interesting calc class. I liked engineering and physics apps as well as volumes

OpenStudy (psymon):

I suck with word problems, haha. I did horribly with some of the mixture ones xD

OpenStudy (psymon):

But maybe thats why we don't have calc 4, we cover more in calc 2.

OpenStudy (raffle_snaffle):

i love application. Did you guys ever talk about exponential growth and decay?

OpenStudy (psymon):

Yep, definitely. Those were np.

OpenStudy (raffle_snaffle):

calc 4 is only vectors.

OpenStudy (psymon):

I see. Lol, the beginning of my calc 3 text is vectors.

OpenStudy (raffle_snaffle):

are you in college or high school?

OpenStudy (psymon):

College. I'm a math tutor where I go to school. I'm just the least experienced one there, lol.

OpenStudy (raffle_snaffle):

Nice, I was considering signing up to be a math tutor at my school but i don't have time. what college do you go to?

OpenStudy (psymon):

I'm in Vegas. I'm going to start my first semester at UNLV on the 26th, so I'll be doing part time at College of Southern Nevada and UNLV. One of the plans in mind, though: 東京でテンプル大学に入るつもりです。 Plan on studying at a university in Tokyo. I don't wanna stay in Vegas, haha.

OpenStudy (raffle_snaffle):

Why Tokyo and what are you studying?

OpenStudy (psymon):

Well, Tokyo is a Japanese university, but it's home school is Temple, which is in Philadelphia. Not only am I studying math, but also Japanese. I wanna leave Vegas and just be able to actually see Japan. I thought it'd be fun to study there and, since its a branch of an American uni, I can still get financial aid and other benefits that U.S students normally get.

OpenStudy (psymon):

Bleh, *in Tokyo

OpenStudy (raffle_snaffle):

Oh very nice. One of my buddies has about a term left until he gets his BS in mathematics. in fact he is on this site. You could pick his brains when you see him.

OpenStudy (psymon):

Right. I'm working on becoming a math teacher, so this site is awesome experience and review.

OpenStudy (raffle_snaffle):

he knows a crap load of math. He actually tutors me at my school, very smart guy

OpenStudy (raffle_snaffle):

have you taken any analysis math courses or abstract algebra courses?

OpenStudy (psymon):

A lot of the people on this site are geniuses, lol. Even some of the people we help are geniuses considering we're helping them woith calc 3 and Diff EQs material.

OpenStudy (psymon):

I'm about to. Going to take Discrete Math along with Diff EQs in 10 days xD

OpenStudy (raffle_snaffle):

i won't rake diff eqs until next spring.

OpenStudy (psymon):

And I wont offically take calc 3 until spring, but I have the text and am studying. So if you're going to do the same material then that's great. I'm looking at a section about determining the distance from a plane to a point not on theplane.

OpenStudy (raffle_snaffle):

are you going to be learning about sequence and series in calc 3?

OpenStudy (psymon):

Did all of that in calc 2 already.

OpenStudy (raffle_snaffle):

I see.. lol well i didn't like series. Power and tayler series isn't too bad though

OpenStudy (psymon):

Some of them get tricky for sure, lol. I just need practice on them.

OpenStudy (raffle_snaffle):

well, I am going to solve this vector problem, appreciate the help and good luck to you!

OpenStudy (psymon):

Alright, let us know if you need the help :3

OpenStudy (raffle_snaffle):

so i got <-1/6, -1/6, 1/3>

OpenStudy (psymon):

Before I even check the answer myself it has to be wrong. It must be a unit vector, which means \[a ^{2}+b ^{2}+c ^{2} = 1\]

OpenStudy (raffle_snaffle):

what do you mean? I manipulated the problem wrong?

OpenStudy (psymon):

Seems so.

OpenStudy (raffle_snaffle):

it's all your fault=P

OpenStudy (psymon):

Hmm.....Okay, my bad, we didn't finish the problem. You still need to check your math on thefirst part, though xD This portion wasn't supposed to give us a unit vector, only the regular vector. But I still got a different answer. The vector I came up with was <-3/14, 2/14, -1/14>

OpenStudy (raffle_snaffle):

what did you get for the mag of a?

OpenStudy (psymon):

We don't use mag a, we use mag b.

OpenStudy (raffle_snaffle):

oh so we are looking for the projection of a onto b?

OpenStudy (psymon):

Right.

OpenStudy (raffle_snaffle):

okay well there is my problem...

OpenStudy (psymon):

Nope, my problem, you were right, didn't notice it said b x_x

OpenStudy (raffle_snaffle):

soooo my math isn't wrong?

OpenStudy (psymon):

Nope, just checked it.

OpenStudy (psymon):

I just chose the wrong projection.

OpenStudy (psymon):

Alright, so now we just need to divide this vector by it's magnitude to get the unit vector.

OpenStudy (raffle_snaffle):

mag of a, right?

OpenStudy (psymon):

magnitude of the resulting vector. So the 1/6 vector you had.

OpenStudy (raffle_snaffle):

shoot, i just erased all my work. lol

OpenStudy (psymon):

Okay, got the answer, just needed to check it x_x

OpenStudy (psymon):

yeah, we have the correct vector, which is the projection you came up with, but if we want the unit vector version of it, we need to divide it by its own magnitude.

OpenStudy (raffle_snaffle):

for the first one i get -sqrt14/84

OpenStudy (psymon):

For thei component?

OpenStudy (raffle_snaffle):

wait by vector b or vector a mag?

OpenStudy (raffle_snaffle):

yeh for the i comp

OpenStudy (psymon):

So we have the vector <-1/6, -1/6, 2/6> So now we divide this vector by it's magnitude: \[\sqrt{(\frac{ -1 }{ 6 })^{2}+(\frac{ 1 }{ 6 })^{2}+(\frac{ 2 }{ 6 })^{2}}\] \[\frac{ \sqrt{6} }{ \sqrt{36} }=\frac{ \sqrt{6} }{ 6 }\] That should be the magnitude you come up with. If that makes sense o.o

OpenStudy (raffle_snaffle):

i got 0

OpenStudy (psymon):

O.o Do you see where I got my work from, though?

OpenStudy (raffle_snaffle):

yeah i see what you did there

OpenStudy (psymon):

Yeah, that would be the process. Then you divide your vector by that magnitude: \[\frac{ proj _{b}a }{ ||proj _{b}a|| }= (<\frac{ -1i }{ 6 },\frac{ -1j }{ 6 },\frac{ 2k }{ 6 }>)*\frac{ 6 }{ \sqrt{6} }\]

OpenStudy (raffle_snaffle):

your are mult?

OpenStudy (psymon):

Well, I flipped the fraction to multiply it.

OpenStudy (psymon):

Instead of sqrt(6)/6 and dividing by that, I just flipped it to multiply.

OpenStudy (raffle_snaffle):

ah yes you did

OpenStudy (psymon):

And then multiplying that through would give you your unit vector :3

OpenStudy (raffle_snaffle):

i still get 0

OpenStudy (psymon):

During the magnitude part?

OpenStudy (raffle_snaffle):

hold on here

OpenStudy (psymon):

lol, kk

OpenStudy (raffle_snaffle):

<-sqrt6/6,-sqrt6/6, sqrt6/6>

OpenStudy (psymon):

Last one is sqrt(6)/3 :P

OpenStudy (raffle_snaffle):

i reduced it when i use the projection formula

OpenStudy (raffle_snaffle):

so ha=P

OpenStudy (psymon):

whatev xD As long as you got the right answer, lol.

OpenStudy (raffle_snaffle):

i don't have a solution yet. these are on a pre-exam handout

OpenStudy (psymon):

Ah. Well I just didn trationalize the denominator on mine. Depending on the professor there's no need to.

OpenStudy (raffle_snaffle):

kk

OpenStudy (psymon):

Mhm. When we don't mess up hopefully it makes sense xD

OpenStudy (raffle_snaffle):

okay cool. Thanks again for walking me through the problem.

OpenStudy (psymon):

yep, np. And im still dealing with vectors so maybe I can help out som emore xD Gotta pack up, though, Ill be back soon.

OpenStudy (raffle_snaffle):

yeah okay, later